OpenStudy (anonymous):

Simplify the following expression. 3y – 2x + z – 2x + 3(x + y) A. –x + 4y + z B. –x + 6y + z C. x + 4y + z D. 3x + 4y + z

OpenStudy (anonymous):

\[\LARGE 3y-2x+z-2x+3x+3y\] can you continue ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i can't do it :( i don't get it?

OpenStudy (callisto):

Group the like terms,ie terms with same variable For instance, you can see there are 3 terms with 'x' -> -2x-2x+3x Simplify it, then you'll get (-2-2+3)x = -x Can you do the rest now?

OpenStudy (anonymous):

3y−2x+z−2x+3x+3y (3y+3y)=6y (-2x-2x+3x)=-x z Put it all back together 6y-x+z

OpenStudy (anonymous):

You break it into parts